OpenStudy (anonymous):

integrate 5x^3/(x^2+9)^(1/2) using trig sub

OpenStudy (anonymous):

x=3tan(theta)

OpenStudy (anonymous):

dx=3sec^2(theta)

OpenStudy (anonymous):

seems good so far, if that's not doing the trick, you might want to try a hyperbolic trigonometric substitution.

OpenStudy (psymon):

Now just tack on that sqrt(x^2+9) = 3sec(theta)

OpenStudy (anonymous):

\[5\int\limits_{}^{}\frac{ (3\tan(\theta)(3\sec ^{3}(\theta) }{3\sec(\theta) }\]

OpenStudy (anonymous):

\[45\int\limits_{}^{}\tan(\theta)\sec ^{2}(\theta)\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

I just realized one of my mistakes

OpenStudy (psymon):

Looks like youre answering your own question xD

OpenStudy (anonymous):

haha, it should only be 15 since two of the 3's cancel

OpenStudy (psymon):

Didnt check, you were doing too well on your own, lol.

OpenStudy (anonymous):

So I have \[15\int\limits_{}^{}\sec(\theta)\tan(\theta)\sec(\theta)\]

OpenStudy (anonymous):

\[u=\sec(\theta) du=\sec(\theta)\tan(\theta)\]

OpenStudy (anonymous):

\[15\int\limits_{}^{}u du\]

OpenStudy (anonymous):

\[(\frac{ 15 }{ 1 })(\frac{ 1 }{ 2 })u ^{2}\]

OpenStudy (anonymous):

\[(\frac{ 15 }{ 2 })\sec ^{2}(\theta)\]

OpenStudy (anonymous):

\[\sec(\theta) = \frac{ \sqrt{x ^{2}+9} }{ 3 }\]

OpenStudy (anonymous):

\[\frac{ 15\sqrt{x ^{2}+9} }{ 18 }\]

OpenStudy (anonymous):

Did I do something wrong? It says I'm getting the wrong answer

OpenStudy (psymon):

You didnt square the square root portion.

OpenStudy (anonymous):

ohhhh, lol

OpenStudy (anonymous):

I just eliminated the square root on top and it's still counting it wrong

OpenStudy (psymon):

Alright, ill work it out then.

OpenStudy (anonymous):

any luck?

OpenStudy (psymon):

Yeah, I got it now, im just being dumb, lol x_x

OpenStudy (anonymous):

haha, nice

OpenStudy (psymon):

Shouldve had this after your substitutions;